OpenStudy (anonymous):

What is the product of (sqrt)3 (7b(sqrt)64 - 5b)? Simplify if possible.

OpenStudy (anonymous):

u mean? \[\sqrt{3} * 7b \sqrt{64}-5b\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then it is \[56\sqrt{3}b-5b\]